nightshade0773:

Kylie is studying the change in the water level of Sweetgum Pond as part of a summer science experiment. In the first week of June, the change in the water level was – 4 centimeters. The next week it stormed almost every day, so the water level rose 7 centimeters. What was the change in the water level over the first two weeks of June? it needs to be in centimeters

AZ:

So the first week, the change in the level of water was -4 inches In the second week, the water rose 7 inches So what was the overall change? You just have to add -4 + 7 = ?? and that's the same thing, by the way, as 7 - 4 if you just rearrange it -4 + 7 = 7 + (-4) which is equal to 7 - 4 so what is 7 - 4 = ?